OpenStudy (anonymous):

a car wheel has a 14 inches radius. through what angle does the wheel turn when the car rolls forward 2 ft????

OpenStudy (anonymous):

answer is 98.2 degrees

OpenStudy (anonymous):

can u show me step by dtep:)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

arc length = rad * angle 2 ft = 14 inch * angle 2*12 inch = 14 inch * angle angle = 24/14 inch = 12/7 radians this angle is in radians we need to convert it into degress so for conversion, (12/7)*(180/pi) solve it and u get 1080/11 deg or 98.2 deg

OpenStudy (anonymous):

hope this is understandable and it helped u

OpenStudy (anonymous):

i got it now.... thanks a lot..:)

OpenStudy (anonymous):

can i ask u about this??? how do u know the answer it'll be a radians or a degrees???

hero (hero):

If they give you a length to find, most likely, the answer is going to be in radians.

OpenStudy (anonymous):

bcoz its a standardnotation that if length and arc is in inches, then angle is in radians angle should always be represented in degrees, so need the conversion